OpenStudy (gabylovesyou):

Jannette can repair a refrigerator in 60 minutes. Liza can repair a refrigerator 20 minutes faster. If they work together, how many minutes does it take them to repair the refrigerator? 120 minutes 40 minutes 24 minutes 20 minutes

OpenStudy (gabylovesyou):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`Jannette can repair a refrigerator in 60 minutes. Liza can repair a refrigerator 20 minutes faster` question: how long does it take Liza to repair the fridge if she worked on her own (with no help) ?

jimthompson5910 (jim_thompson5910):

any ideas @Gabylovesyou ?

OpenStudy (gabylovesyou):

60 - 20 ?

jimthompson5910 (jim_thompson5910):

which is 40 min, yes

jimthompson5910 (jim_thompson5910):

x = time it takes Jannette to do the job alone (with no help) y = time it takes Liza to do the job alone (with no help) z = time it takes to do the job if the two worked together All times (x,y,z) are in minutes In this case, x = 60 min y = 40 min z = unknown \[\Large \frac{1}{x} + \frac{1}{y} = \frac{1}{z}\] \[\Large \frac{1}{60} + \frac{1}{40} = \frac{1}{z}\] solve for z

OpenStudy (gabylovesyou):

24 ? sorrry wifi is slow...

jimthompson5910 (jim_thompson5910):

yes, z = 24